Dividing Partnership Income Tyler Hawes and Piper Albright formed a partnership, investing $116,000 and $174,000, respectively....

  1. Dividing Partnership Income

    Tyler Hawes and Piper Albright formed a partnership, investing $116,000 and $174,000, respectively.

    Determine their participation in the year's net income of $280,000 under each of the following independent assumptions:

    1. No agreement concerning division of net income.
    2. Divided in the ratio of original capital investment.
    3. Interest at the rate of 6% allowed on original investments and the remainder divided in the ratio of 2:3.
    4. Salary allowances of $40,000 and $46,000, respectively, and the balance divided equally.
    5. Allowance of interest at the rate of 6% on original investments, salary allowances of $40,000 and $46,000, respectively, and the remainder divided equally.
    Hawes Albright
    (a) $ $
    (b) $ $
    (c) $ $
    (d) $ $
    (e) $ $
Answer #1

B3 A B C D E | ovou AwN (a) (b) (c) (d) (e) Hawes $ 140,000 $ 112,000 $ 112,000 $ 137,000 $ 135,260 Albright $ 140,000 $ 168,

for formulas and calculations, refer to the image below -

А в D Hawes Albright (a) |=280000/2 = 280000/2 |(b) =280000* 116000/(116000+174000) = 280000*174000/(116000+174000) (c) =1160

In case you have any query, kindly ask in comments.